OpenStudy (anonymous):

Matthew earns $8.50/h working at a grocery store. He receives a raise of 15% per hour. Which expression represents how much money Matthew will earn per hour after the raise? A. $8.50 x 3/20 B. $8.50 x 17/20 C. $8.50 x 23/20 D. $8.50 x37/20

OpenStudy (anonymous):

@cjackson2002 @PaulaLovesSchool13 @perl @KittyT @CausticSyndicalist @angel12310

OpenStudy (anonymous):

A.

OpenStudy (anonymous):

thanks Buddy

OpenStudy (anonymous):

15%=3/20 so its a.8.50*3/20

OpenStudy (anonymous):

can you help on a few more

OpenStudy (anonymous):

sure:)

OpenStudy (anonymous):

welcome buddy:)

OpenStudy (anonymous):

Valerie used to spend 10 h/wk watching TV. Now she spends 1/4 less time per week watching TV. Which expressions represent how many hours per week Valerie spends watching TV now? Choose all answers that are correct. A. 10 - (1/4 x 10) B. 10 + (1/4 x 10) C. 3/4 x 10 D. 1/4 x 10

OpenStudy (anonymous):

A and C

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

During his first year of life, Avery gained 12 3/4 lb. During his second year of life, Avery gained 40% less weight than he did during his first year of life. How much weight did Avery gain during his second year of life? Express your answer as a mixed number in simplest form.

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

7 13/20

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

Dustin is mixing concrete. A formula for concrete calls for 3/8 gal of water. Dustin wants to make 7/9 more concrete than the formula creates. How much water should Dustin use? Express your answer as a fraction in simplest form.

OpenStudy (anonymous):

I was right

OpenStudy (anonymous):

Ya ik i got mixed up with 2 q

OpenStudy (anonymous):

this one is the last one

OpenStudy (anonymous):

7/24

OpenStudy (anonymous):

I got 40%

OpenStudy (anonymous):

sorry:(

OpenStudy (anonymous):

u in k12?

OpenStudy (anonymous):

u would multiply 7/9 and 3/8=7/24

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

so am i

OpenStudy (anonymous):

TNVA

OpenStudy (anonymous):

Sorry ask @iGreen

OpenStudy (igreen):

Um..wait what's your question?

OpenStudy (igreen):

@Wolves_Rock2002 @cjackson2002